# Deep Self-Supervised Disturbance Mapping with the OPERA Sentinel-1 Radiometric Terrain Corrected SAR Backscatter Product

### v1

@@ I Introduction
The Earth’s surface is continually changing due to a combination of natural [1, 2, 3] and human-induced factors [4, 5]. While some changes occur gradually over time (e.g., mountain formation, glacial isostatic adjustment), others are abrupt and can have immediate impacts on the environment (e.g., landslides, wildfires, flooding, deforestation, mining). ect and monitor these surface changes and disturbances over large areas and on various time scales to address critical environmental and societal challenges [6, 7]. Through disturbance delineation, we enhance our capability to monitor vulnerable or protected habitats [6], map the impacts of human urbanization and development [8], track population migration patterns [9], identify tropical forest loss [5, 10], and respond to natural hazards such as earthquakes [11], floods [12, 13], landslides [14, 15], and fires [16, 17].
Among remote sensing techniques, synthetic aperture radar (SAR) is an invaluable modality for disturbance mapping, with the capability to rapidly image the ground regardless of cloud cover or illumination conditions. While optical imagery provides information about spectral changes, SAR provides information about structural changes that are captured via backscatter and phase. Applications include centimeter-scale motion of the Earth’s surface [18, 19], structural damages after earthquakes [11], and landslide [20, 21], and wildfire delineation [17, 16].
@@
Supervised machine learning techniques rely on the existence of a large, labeled training set. However, labeling such datasets for disturbance mapping - particularly global-scale SAR time series spanning months or years - is impractical: it requires expertise of SAR physical mechanisms [25, 26] as well as knowledge of the land cover being monitored to determine what signals are phenological and which are true disturbances. Despite this, many previous works on SAR disturbance mapping are supervised [16, 27, 28, 29], requiring carefully curated, labeled training sets. Not only is this difficult to scale to larger datasets, but the models trained on these specialized datasets lack the generality to be a broadly applicable, generic disturbance model. While there are numerous optical disturbance datasets [30, 31, 32, 33], training a SAR model on optically-derived labels presents significant challenges as some disturbances visible to an optical sensor will be invisible to SAR and vice versa. Additionally, with forthcoming NASA ISRO SAR (NISAR) data, where the creation of a large labeled corpus from the new L-band imagery is not possible, the demonstration of a label-free approach for SAR-based applications is essential.
@@ I-A Our Contributions
In this work, we train a self-supervised vision transformer [41] to estimate the per-pixel distribution from baseline of OPERA RTC-S1 imagery. This estimated distribution can be compared to a new acquisition by computing the Mahalanobis distance, yielding a disturbance metric that can be used to delineate disturbance. The metric has a statistical interpretation as the number of standard deviations (SD) from the modeled per-pixel mean. More simply put, the metric quantifies how likely a pixel is disturbed, with larger values indicating greater likelihood of disturbance. The vision transformer is trained in a self-supervised manner, meaning no labels or human annotation are required; the only input is the OPERA RTC-S1 imagery. We evaluate our disturbance metric on external validation data from three recent natural disasters: 1) a landslide in Papua New Guinea, 2) a series of fires in Chile, and 3) flooding in Bangladesh. We use the same model and metric for every event. The validation data is derived from publicly-available damage maps [44, 45]; much of which is made from Very High Resolution (VHR) optically-derived inputs. These natural disasters represent high-intensity disturbances where we expect significant alignment between optical- and SAR-based disturbance delineations. The validation data used for this work, including the disturbance delineations and data provenance, are found in the GitHub repository [46]. We demonstrate that the proposed vision transformer-based disturbance model has systematically better performance compared to an RNN-based model [11] and the classical log ratio method for SAR change mapping [17]. Code for our data [^2] https://github.com/OPERA-Cal-Val/dist-s1-events and experiments [^3] https://github.com/opera-adt/distmetrics are available via GitHub.
@@ II-A Disturbance Mapping
Currently, there are two operational [^5] Here, operational here means that each new SAR acquisition triggers the generation of a new disturbance map. SAR disturbance products, which track tropical forest loss: ALOS forest alerts [51, 52] and RAdar for Detecting Deforestation (RADD) [10, 53]. Both rely on estimating backscatter conditioned on “forest” and “non-forest” distributions and delineating disturbance when there is evidence forest has been removed based on these modeled distributions. Our approach aims to delineate a broader class of disturbances, across terrains and disturbance events. To this end, we utilize a vision transformer trained on near-global RTC-S1 data to determine the range of nominal backscatter values in varied environments and land cover types, not just dense tropical forest.
@@
Figure 2: A flow chart illustrating the transformer metric proposed in this paper. The metric is computed for dual polarization images (see Figure 1 for 1-dimensional visualization). The first arrow is where the deep-learning model is utilized to estimate the per-pixel distribution of the $x_{T+1}$.
@@
Above, $I_{1}$ is the post-event cross- or co-polarization image and $I_{0}$ is a pre-event cross- or co-polarization image. $I_{0}$ is also sometimes an aggregated reference image; in our experiments, we will set $I_{0}$ to be the median of the available pre-event imagery. It is well known that large negative values in $\log(I_{1})-\log(I_{0})$ indicate vegetation loss when we consider cross-polarization data [17, 25]. Above, we consider the absolute value because we aim to delineate generic disturbances from a baseline set of imagery, which could be indicated by anomalously low or high measurements in either polarization. In Section III-B, we will define a disturbance metric for our transformer-based approach, which compares favorably to $\ell$ in our experiments (Section VI).
@@ III-C Model Architecture, Hyperparameters, and Training
Our transformer encoder uses a model dimension of 256, 4 attention heads, 4 layers, and a feedforward dimension of 768. The input to the model is a sequence of $T=$ 2 to 10 SAR images. In each batch, $T$ is chosen randomly in this range so the model can learn to forecast the distributions with varying amounts of pre-event information. Operationally, this is important as data may often be missing or an area of interest lacks a long acquisition history. For example, Sentinel-1 data does not consistently image in 12-day increments, so in some cases many months can go by before additional data is collected. Our validation datasets discussed in Section VI have 10, 7, and 4 pre-event images available, underscoring the need for a temporally flexible model.
@@ IV RTC-S1 Data
The OPERA RTC-S1 provides measurements of the $\gamma^{0}$ radar backscatter corrected for radiometry and topography [23, 24]. All RTC-S1 products can be accessed through the Alaska Satellite Facility Distributed Active Archive Center (ASF DAAC) [96]. These corrections result in an analysis-ready product that primarily captures signals related to the physical characteristics of ground-scattering surfaces, such as surface roughness, soil moisture, and vegetation structure. The RTC-S1 is generated from Copernicus Sentinel-1 (S1) interferometric wide (IW) single-look complex (SLC) data, provided by the European Space Agency (ESA). Sentinel-1 constellation can currently acquire data at 12 day repeat pass temporal frequency, i.e. with respect to a fixed orbit track. The ground sampling when independent of orbit track is increased to between 1 and 6 days; this combines ascending (satellite flying north and looking east) and descending (satellite flying south and looking west) geometries. The imaging mode and temporal coverage of Sentinel-1 data varies with location on Earth [49]. The RTC-S1 product has a near-global geographic scope that covers all landmasses, excluding Antarctica, and its temporal frequency aligns with the availability of Sentinel-1A SLC data. RTC-S1 currently only includes data from Sentinel-1A satellite because OPERA started forward mode production in October 2023 after the loss of Sentinel-1B. Each RTC-S1 product corresponds to a single Sentinel-1 burst and is mapped onto standardized Universal Transverse Mercator (UTM) or Polar Stereographic projection systems.
@@ V Disaster Data
Model Landslide Fire Flood
PR AUC Max $F_{1}$ Score PR AUC Max $F_{1}$ Score PR AUC Max $F_{1}$ Score
@@
TABLE I: Summary of our results. We report Area Under the Precision-Recall curve (PR AUC), which is an overall measurement of classifier quality, and the maximum $F_{1}$ score obtained by each model. Best scores are bolded. Across all validation datasets and metrics, the transformer performs the best.
@@
On May 24, 2024, a large catastrophic landslide occurred near Yambali, Enga Province, Papua New Guinea (see Figure 3) [103, 104]. News media reported that that the landslide buried numerous villages, with the death toll estimated to be between 650 and 2,000 people [103]; the official death toll at time of writing is 670 [105]. In addition, approximately 7,000 people have been displaced by the event [103].
Our input to the model is a sequence of ten pre-event OPERA RTC-S1 SAR images from 1/23/24 to 5/22/24 at 12 day intervals (except 5/10/204 when there was no data distributed despite a Sentinel-1 pass) and a post-event image from 6/3/2024. To delineate disturbances, we compute the Mahalanobis distance between the models’ estimate distribution and this post-event image as described in Section III-B. We analyze an area of approximately 6 km $\times$ 6 km. We manually mapped the landslide extent using PlanetScope optical imagery (3 m pixel) [89], see Figure 5.
@@
In early February of 2024, a series of wildfires began in the Valparaíso region of Chile [106]. Exacerbated by the historic drought in the region and the 2023 El Niño weather event, by February 3 the fires had burned over 8,500 hectares [106]. The death toll is at least 131 people [107]. Figure 4 shows Landsat 8 and 9 false color imagery of two burn scars[90].
For this site, The Copernicus Emergency Management Service (CEMS) [108] produced a disturbance map of the region that we use as our ground truth, see Figure 6. The area of interest is approximately 45 km $\times$ 33 km. The data was generated by a variety of VHR sources and was continually updated for several months after the fires started [109]. Our RTC-S1 data used included all seven available pre-event SAR images from 11/12/24 to 1/23/24 (all Sentinel-1 passes resulted in distributed data) and one post-event SAR image from 2/29/24. Although there was data collected and distributed on February 5 and 17 by Sentinel-1 as the fires were ongoing, there was noticeable increase in disturbance in the February 29 data. This acquisition was also closer in time to the VHR map, which was updated well beyond the fire’s start [109].
@@
On May 26, 2024, Cyclone Remal landed in Bangladesh, causing major flooding that has continued throughout the summer [110]. At least 4.6 million people were affected throughout the summer [110]. Because the flooding occurred throughout the summer, it was challenging to establish a baseline of imagery in the RTC-S1 archive in which these floods were not present. We selected validation data containing 4 pre-event images from entirely before the cyclone, acquired between 3/31/24 and 5/18/24. We use a ground truth damage map from the United Nations Satellite Center (UNOSAT) [44] which was derived from a Sentinel-1 image on July 4th [102]. We use the nearest RTC-S1 post-event image, which was acquired July 5th, 2024. Our flooding analysis is a subset of the country along Ganges delta around the Baleshwari River. The area of interest is approximately 42 km $\times$ 30 km, see Figure 7.
@@ VI-A Evaluation Procedure and Performance Metrics
Figure 8: The transformer metric, the RNN metric, and the log-ratio metric (see Section III-B) on the Papua New Guinea scene before the landslide (left column) and after the landslide (right column). A strong model should produce a metric with low values in the left column where we expect little or no disturbance since no landslide has occurred yet, while also accurately and confidently capturing the landslide extents post-landslide (right column). Note the qualitative strength of the transformer compared to the classical log ratio, which is much noisier and less confident in identifying the landslide disturbance.
(a) Transformer
(b) RNN
(c) Log Ratio
Figure 9: Best (measured by $F_{1}$ score) binary damage maps from each method. The two deep methods identify more of the landslide while also having fewer false positives than the log ratio methods.
Our analysis combines qualitative and quantitative results to facilitate visual comparisons and highlight the strength of our approach. Qualitative results include two map types: (1) disturbance metric maps visualizing $d$ (transformer or RNN) or $\ell$ (log ratio) across the area of interest, and (2) binary disturbance delineations based on the optimal threshold $\tau$ (according to $F_{1}$ score, see below). Operationally, ground truth information is not available, so these maps are the critical outputs that delineate disturbances. In our experiments, these maps highlight the transformer’s effectiveness in identifying disturbances while limiting false positives.
@@
Figure 8 shows the transformer metric (see Section III-B), the RNN metric, and the log ratio metric. The first column is the final pre-event image and the second column is the post-event image (see Section V-A). The two deep methods, particularly the transformer, are significantly more accurate than the log ratio with respect to our validation data, identifying the landslide extents more confidently and limiting the number of false positives particularly in the first pre-event image.
Figure 10 shows the PR curves for the landslide scenes. The transformer achieves by far the best AUC ($0.732$) indicating overall model quality independent of threshold $\tau$. The RNN is second with an AUC of $0.642$, and both deep methods significantly outperform the classical log ratio method (AUC of $0.067$), which barely outperforms random guessing. The optimal $\tau$ and corresponding $F_{1}$ score is starred. The transformer obtains the best $F_{1}$ score of $0.769$, the RNN $0.699$, and the log ratio $0.194$. The disturbance delineations corresponding to the optimal $\tau$ are shown in Figure 9. While, in practice, the optimal $\tau$ would not be known a priori, we include these figures to highlight an example disturbance delineation from each model.
Figure 10: PR curves for the landslide scene. The transformer achieves the largest area under the curve (AUC). The transformer also obtains the best $F_{1}$ score (starred for each model). The disturbance delineations corresponding to the optimal $\tau$ that obtains optimal $F_{1}$ score are shown in Figure 9.
Figure 11: PR curves for the Chilean fire scene. The transformer is the most performant model, achieving the largest area under the curve. The transformer also obtains the best $F_{1}$ score (starred for each model). The binary damage maps corresponding to these starred $F_{1}$ scores are shown in Figure 13.
@@
Figure 12 shows disturbance metrics for the transformer, RNN, and log-ratio as in the previous section. As before, the transformer is the best performing, followed by the RNN. Both deep methods are significantly more accurate than the log ratio, both in terms of identifying the fire extents confidently and limiting the number of false positives elsewhere in the scene, or before the fire occurred.
@@
Figure 12: The disturbance metrics on the Chilean fire scene before the fire (left column) and after (right column). Note the qualitative strength of the two deep methods - particularly the transformer - compared to the classical log ratio, where there is more high-confidence disturbance away from or before the fire.
@@
Figure 13: Best (measured by $F_{1}$ score) binary damage maps for the Chilean fires from each method. As with the landslide, the two deep methods - particularly the transformer - identify more of the fire extents while also having significantly fewer false positives than the log ratio methods.
Figure 11 shows the PR curves for the Chilean fire scenes. As with the Papua New Guinea landslide, the transformer achieves the best AUC of $0.680$. The RNN is a second with an AUC of $0.624$, and both deep methods significantly outperform the log ratio ($0.391$). The transformer also obtains the best $F_{1}$ score of $0.645$ (the best $F_{1}$ obtained by each method is starred). The damage maps corresponding to those best $F_{1}$ scores are shown in Figure 13. These binary damage maps reflect the deep learning methods’ - especially the transformer’s - ability to accurately map damage extents while limiting false positives.
VI-D 2024 Bangladesh Floods
As in the previous two sites, we study the transformer metric, the RNN metric, and the log-ratio metric for the pre-event and post-event image in Figure 14. The disturbance is more extensive in this scene and it is difficult to draw any immediate qualitative conclusions from these maps. However, one interesting observation is that the log ratio method is somewhat confident that parts of the river are a flooded area; the deep methods do not suffer as much from this (per the ground truth damage map in Figure 7, the river is not considered a flooded area).
As with the previous subsections, the quantitative results show that the transformer is the most performant of the methods. Figure 16 shows the PR curves for the flood scene. Again, the transformer achieves the best AUC $0.754$. The RNN is quite similar to the classical log ratio method (AUCs of $0.705$ and $0.715$, respectively) underscoring the improved performance from the transformer compared to the RNN. The transformer also obtains the best $F_{1}$ score ($0.701$). The disturbance delineations corresponding to those best $F_{1}$ scores are shown in Figure 15.
(a) Transformer
(b) RNN
(c) Log Ratio
Figure 14: Disturbance metrics on the Bangladesh scene before the flood (left column) and after (right column). Note the log ratio methods seem (incorrectly) more confident than the deep methods that the upper part of the Baleshwari River is a flood area.
(a) Transformer
(b) RNN
(c) Log Ratio
Figure 15: Best (measured by $F_{1}$ score) binary disturbance maps for the flooding scene from each method. As with the previous scenes, the two deep methods, particularly the transformer, identify more of the flooding while also having fewer false positives than the log ratio methods.
Figure 16: PR curves for the Bangladesh flood scene. The transformer is the most performant model, achieving the largest area under the curve. The transformer also obtains the best $F_{1}$ score (starred for each model). The binary disturbance maps corresponding to the starred $F_{1}$ scores are shown in Figure 15.

### v2

@@ I Introduction
The Earth’s surface is continually changing due to a combination of natural [1, 2, 3] and human-induced factors [4, 5]. While some changes occur gradually over time (e.g., mountain formation, glacial isostatic adjustment), others are abrupt and can have immediate impacts on the environment (e.g., landslides, wildfires, flooding, deforestation, mining). Satellite remote sensing can detect these surface changes and disturbances over large areas and on various time scales to address critical environmental and societal challenges [6, 7]. Through disturbance delineation, we enhance our capability to monitor vulnerable or protected habitats [6], map the impacts of human urbanization and development [8], track population migration patterns [9], identify tropical forest loss [5, 10], and respond to natural hazards such as earthquakes [11], floods [12, 13], landslides [14, 15], and fires [16, 17].
Among remote sensing techniques, synthetic aperture radar (SAR) is an invaluable modality for disturbance mapping, with the capability to rapidly image the ground regardless of cloud cover or illumination conditions. While optical imagery provides information about spectral changes, SAR provides information about structural changes that are captured via backscatter and phase. Applications include centimeter-scale motion of the Earth’s surface [18, 19], crop mapping [20], structural damages after earthquakes [11], and landslide [21, 22], and wildfire delineation [17, 16].
@@
Supervised machine learning techniques rely on the existence of a large, labeled training set. However, labeling such datasets for disturbance mapping - particularly global-scale SAR time series spanning months or years - is impractical: it requires expertise of SAR physical mechanisms [26, 27] as well as knowledge of the land cover being monitored to determine what signals are phenological and which are true disturbances. Despite this, many previous works on SAR disturbance mapping are supervised [16, 28, 29, 30], requiring carefully curated, labeled training sets. Not only is this difficult to scale to larger datasets, but the models trained on these specialized datasets lack the generality to be a broadly applicable, generic disturbance model. While there are numerous optical disturbance datasets [31, 32, 33, 34] and corresponding supervised models (see e.g. [35]), training a SAR model on optically-derived labels presents significant challenges as some disturbances visible to an optical sensor will be invisible to SAR and vice versa. Additionally, with forthcoming NASA ISRO SAR (NISAR) data, where the creation of a large labeled corpus from the new L-band imagery is not possible, the demonstration of a label-free approach for SAR-based applications is essential.
@@ I-A Our Contributions
In this work, we train a self-supervised vision transformer [43] on near-global OPERA RTC-S1 imagery and demonstrate superior damage mapping performance across environments and disaster events compared to previous works. Our transformer learns to estimate the per-pixel distribution from baseline of OPERA RTC-S1 imagery. This estimated distribution can be compared to a new acquisition by computing the Mahalanobis distance, yielding a disturbance metric that can be used to delineate disturbance. The metric has a statistical interpretation as the number of standard deviations (SD) from the modeled per-pixel mean. More simply put, the metric quantifies how likely a pixel is disturbed, with larger values indicating greater likelihood of disturbance. The vision transformer is trained in a self-supervised manner, meaning no labels or human annotation are required; the only input is the OPERA RTC-S1 imagery. We evaluate our disturbance metric on external validation data from three recent natural disasters: 1) a landslide in Papua New Guinea, 2) a series of fires in Chile, and 3) flooding in Bangladesh. We use the same model and metric for every event. The validation data is derived from publicly-available damage maps [46, 47]; much of which is made from Very High Resolution (VHR) optically-derived inputs. These natural disasters represent high-intensity disturbances where we expect significant alignment between optical- and SAR-based disturbance delineations. The validation data used for this work, including the disturbance delineations and data provenance, are found in the GitHub repository [48]. We demonstrate that the proposed vision transformer-based disturbance model has systematically better performance compared to an RNN-based model [11] and the classical log ratio method for SAR change mapping [17]. Code for our data [^2] https://github.com/OPERA-Cal-Val/dist-s1-events and experiments [^3] https://github.com/opera-adt/distmetrics are available via GitHub.
@@ II-A Disturbance Mapping
Currently, there are two operational [^5] Here, operational here means that each new SAR acquisition triggers the generation of a new disturbance map. SAR disturbance products, which track tropical forest loss: Advanced Land Observing Satellite (ALOS) forest alerts [53, 54] and RAdar for Detecting Deforestation (RADD) [10, 55]. Both rely on estimating backscatter conditioned on “forest” and “non-forest” distributions and delineating disturbance when there is evidence forest has been removed based on these modeled distributions. Our approach aims to delineate a broader class of disturbances, across terrains and disturbance events. To this end, we utilize a vision transformer trained on near-global RTC-S1 data to determine the range of nominal backscatter values in varied environments and land cover types, not just dense tropical forest.
@@
Figure 2: A flow chart illustrating the transformer metric proposed in this paper. The metric is computed for dual polarization images (see Fig. 1 for 1-dimensional visualization). The first arrow is where the deep-learning model is utilized to estimate the per-pixel distribution of the $x_{T+1}$.
@@
Above, $I_{1}$ is the post-event cross- or co-polarization image and $I_{0}$ is a pre-event cross- or co-polarization image. $I_{0}$ is also sometimes an aggregated reference image; in our experiments, we will set $I_{0}$ to be the median of the available pre-event imagery. Although it was developed over 30 years ago, log ratio is still widely used in modern SAR damage mapping research [57, 14, 21, 58, 59, 60, 61]. It is well known that large negative values in $\log(I_{1})-\log(I_{0})$ indicate vegetation loss when we consider cross-polarization data [17, 26]. Above, we consider the absolute value because we aim to delineate generic disturbances from a baseline set of imagery, which could be indicated by anomalously low or high measurements in either polarization. In Section III-B, we will define a disturbance metric for our transformer-based approach, which compares favorably to $\ell$ in our experiments (Section VI).
@@ III-C Model Architecture, Hyperparameters, and Training
Our transformer encoder uses a model dimension of 256, 4 attention heads, 4 layers, and a feedforward dimension of 768. The input to the model is a sequence of $T$ SAR images, each with size $2\times 16\times 16$ (channels, heigh, and width, respectively, for a total input size of $T\times 2\times 16\times 16$. The channel dimension $2$ is due to the dual-polarization SAR imagery (VV/VH). In each batch, $T$ is chosen randomly to be between 2 and 10 so the model can learn to forecast the distributions with varying amounts of pre-event information. Operationally, this is important as data may often be missing or an area of interest lacks a long acquisition history. For example, Sentinel-1 data does not consistently image in 12-day increments, so in some cases many months can go by before additional data is collected. Our validation datasets discussed in Section VI have 10, 7, and 4 pre-event images available, underscoring the need for a temporally flexible model. Future work will focus on scaling up the model to allow for larger values of $T$.
@@ IV RTC-S1 Data
The OPERA RTC-S1 provides measurements of the radar backscatter, $\gamma^{0}$, corrected for radiometry and topography [24, 25]. All RTC-S1 products can be accessed through the Alaska Satellite Facility Distributed Active Archive Center (ASF DAAC) [103]. These corrections result in an analysis-ready product that primarily captures signals related to the physical characteristics of ground-scattering surfaces, such as surface roughness, soil moisture, and vegetation structure. The RTC-S1 is generated from Copernicus Sentinel-1 (S1) interferometric wide (IW) single-look complex (SLC) data, provided by the European Space Agency (ESA). Sentinel-1 constellation can currently acquire data at 12 day repeat pass temporal frequency, i.e. with respect to a fixed orbit track. The ground sampling when independent of orbit track is increased to between 1 and 6 days; this combines ascending (satellite flying north and looking east) and descending (satellite flying south and looking west) geometries. The imaging mode and temporal coverage of Sentinel-1 data varies with location on Earth [51]. The RTC-S1 product has a near-global geographic scope that covers all landmasses, excluding Antarctica, and its temporal frequency aligns with the availability of Sentinel-1A SLC data. RTC-S1 currently only includes data from Sentinel-1A satellite because OPERA started forward mode production in October 2023 after the loss of Sentinel-1B. Each RTC-S1 product corresponds to a single Sentinel-1 burst and is mapped onto standardized Universal Transverse Mercator (UTM) or Polar Stereographic projection systems.
@@ V Disaster Data
TABLE I: Summary of Results
Landslide Fire Flood
Model PR AUC Max $F_{1}$ Score PR AUC Max $F_{1}$ Score PR AUC Max $F_{1}$ Score
@@
On May 24, 2024, a large catastrophic landslide occurred near Yambali, Enga Province, Papua New Guinea (see Fig. 3) [110, 111]. News media reported that that the landslide buried numerous villages, with the death toll estimated to be between 650 and 2,000 people [110]; the official death toll at time of writing is 670 [112]. In addition, approximately 7,000 people have been displaced by the event [110].
Our input to the model is a sequence of ten pre-event OPERA RTC-S1 SAR images from 1/23/24 to 5/22/24 at 12 day intervals (except 5/10/24 when there was no data distributed despite a Sentinel-1 pass) and a post-event image from 6/3/2024. To delineate disturbances, we compute the Mahalanobis distance between the models’ estimate distribution and this post-event image as described in Section III-B. We analyze an area of approximately 6 km $\times$ 6 km. We manually mapped the landslide extent using PlanetScope optical imagery (3 m pixel) [96], see Fig. 5.
@@
In early February of 2024, a series of wildfires began in the Valparaíso region of Chile [113]. Exacerbated by the historic drought in the region and the 2023 El Niño weather event, by February 3 the fires had burned over 8,500 hectares [113]. The death toll is at least 131 people [114]. Fig. 4 shows Landsat 8 and 9 false color imagery of two burn scars[97].
For this site, The Copernicus Emergency Management Service (CEMS) [115] produced a disturbance map of the region that we use as our ground truth, see Fig. 6. The area of interest is approximately 45 km $\times$ 33 km. The data was generated by a variety of VHR sources and was continuously updated for several months after the fires started [116]. Our RTC-S1 data used included all seven available pre-event SAR images from 11/12/24 to 1/23/24 (all Sentinel-1 passes resulted in distributed data) and one post-event SAR image from 2/29/24. Although there was data collected and distributed on February 5 and 17 by Sentinel-1 as the fires were ongoing, there was noticeable increase in disturbance in the February 29 data. This acquisition was also closer in time to the VHR map, which was updated well beyond the fire’s start [116].
@@
On May 26, 2024, Cyclone Remal landed in Bangladesh, causing major flooding that has continued throughout the summer [117]. At least 4.6 million people were affected throughout the summer [117]. Because the flooding occurred throughout the summer, it was challenging to establish a baseline of imagery in the RTC-S1 archive in which these floods were not present. We selected validation data containing 4 pre-event images from entirely before the cyclone, acquired between 3/31/24 and 5/18/24. We use a ground truth damage map from the United Nations Satellite Center (UNOSAT) [46] which was derived from a Sentinel-1 image on July 4th [109]. We use the nearest RTC-S1 post-event image, which was acquired July 5th, 2024. Our flooding analysis is a subset of the country along Ganges delta around the Baleshwari River. The area of interest is approximately 42 km $\times$ 30 km, see Fig. 7.
@@ VI-A Evaluation Procedure and Performance Metrics
Figure 8: Disturbance metrics, damage maps, and ground truth data from the landslide in Papua New Guinea. Each row corresponds to a different model. The first two columns show the model’s disturbance metric (see Section III-B) before the landslide (first column) and after the landslide (second column). A strong model should produce a metric with low values pre-landslide (when we expect little or no disturbance since no landslide has occurred yet) while also accurately and confidently capturing the damage extents post-landslide. The third and fourth columns show the best (in terms of $F_{1}$ score, see Fig. 9) binary damage maps derived from each method’s metric. The rightmost column shows the ground truth damage map, reproduced from Fig. 5 (right panel) for convenience. The transformer metric is the strongest, resulting in the damage map with the fewest false positives while also accurately capturing the extents of the landslide.
Our analysis combines qualitative and quantitative results to facilitate visual comparisons and highlight the strength of our approach. Qualitative results include two map types: (1) disturbance metric maps visualizing $d$ (transformer or RNN) or $\ell$ (log ratio) across the area of interest, and (2) binary (i.e., disturbance or no disturbance) disturbance delineations based on the optimal threshold $\tau$ (according to $F_{1}$ score, see below). Operationally, ground truth information is not available, so these maps are the critical outputs that delineate disturbances. In our experiments, these maps highlight the transformer’s effectiveness in identifying disturbances while limiting false positives.
@@
Fig. 8 shows the transformer metric (see Section III-B), the RNN metric, and the log ratio metric. The first column is the final pre-event image and the second column is the post-event image (see Section V-A). The two deep methods, particularly the transformer, are significantly more accurate than the log ratio with respect to our validation data, identifying the landslide extents more confidently and limiting the number of false positives particularly in the first pre-event image.
Fig. 9 shows the PR curves for the landslide scenes. The transformer achieves by far the best AUC ($0.732$) indicating overall model quality independent of threshold $\tau$. The RNN is second with an AUC of $0.642$, and both deep methods significantly outperform the classical log ratio method (AUC of $0.067$), which barely outperforms random guessing. The optimal $\tau$ and corresponding $F_{1}$ score is starred. The transformer obtains the best $F_{1}$ score of $0.769$, the RNN $0.699$, and the log ratio $0.194$. The disturbance delineations corresponding to the optimal $\tau$ are shown in Fig. 8, third and fourth columns. While in practice, the optimal $\tau$ would not be known a priori, we include these figures to highlight an example disturbance delineation from each model.
Figure 9: PR curves for the landslide scene. The transformer achieves the largest area under the curve (AUC). The transformer also obtains the best $F_{1}$ score (starred for each model). The disturbance delineations corresponding to the optimal $\tau$ that obtains optimal $F_{1}$ score are shown in Fig. 8, third and fourth columns.
Figure 10: PR curves for the Chilean fire scene. The transformer is the most performant model, achieving the largest area under the curve. The transformer also obtains the best $F_{1}$ score (starred for each model). The binary damage maps corresponding to these starred $F_{1}$ scores are shown in Fig. 11, third and fourth columns.
@@
Fig. 11 shows disturbance metrics for the transformer, RNN, and log-ratio as in the previous section. As before, the transformer is the best performing, followed by the RNN. Both deep methods are significantly more accurate than the log ratio, both in terms of identifying the fire extents confidently and limiting the number of false positives elsewhere in the scene, or before the fire occurred.
@@
Figure 11: Disturbance metrics, damage maps, and ground truth data from the fire in Chile. Each row corresponds to a different model. The first two columns show the model’s disturbance metric (see Section III-B) before the fire (first column) and after the fire (second column). The third and fourth columns show the best (in terms of $F_{1}$ score, see Fig. 10) binary damage maps derived from each method’s metric. The rightmost column shows the ground truth damage map, reproduced from Fig. 6 (right panel) for convenience. As with the landslide, the transformer metric most accurately and confidently captures the fire extents while limiting false positives elsewhere.
Fig. 10 shows the PR curves for the Chilean fire scenes. As with the Papua New Guinea landslide, the transformer achieves the best AUC of $0.680$. The RNN is a second with an AUC of $0.624$, and both deep methods significantly outperform the log ratio ($0.391$). The transformer also obtains the best $F_{1}$ score of $0.645$ (the best $F_{1}$ obtained by each method is starred). The damage maps corresponding to those best $F_{1}$ scores are shown in Fig. 11, third and fourth columns. These binary damage maps reflect the deep learning methods’ - especially the transformer’s - ability to accurately map damage extents while limiting false positives.
VI-D 2024 Bangladesh Floods
As in the previous two sites, we study the transformer metric, the RNN metric, and the log-ratio metric for the pre-event and post-event image in Fig. 12. The disturbance is more extensive in this scene and it is difficult to draw any immediate qualitative conclusions from these maps. However, one interesting observation is that the log ratio method is somewhat confident that parts of the river are a flooded area; the deep methods do not suffer as much from this (per the ground truth damage map in Fig. 7, the river is not considered a flooded area).
As with the previous subsections, the quantitative results show that the transformer is the most performant of the methods. Fig. 13 shows the PR curves for the flood scene. Again, the transformer achieves the best AUC $0.754$. The RNN is quite similar to the classical log ratio method (AUCs of $0.705$ and $0.715$, respectively) underscoring the improved performance from the transformer compared to the RNN. The transformer also obtains the best $F_{1}$ score ($0.701$). The disturbance delineations corresponding to those best $F_{1}$ scores are shown in Fig. 12, third and fourth columns.
@@
Figure 12: Disturbance metrics, damage maps, and ground truth data from the flooding in Bangladesh. Each row corresponds to a different model. The first two columns show the model’s disturbance metric (see Section III-B) before the flood (first column) and after (second column). The third and fourth columns show the best (in terms of $F_{1}$ score, see Fig. 13) binary damage maps derived from each method’s metric. The rightmost column shows the ground truth damage map, reproduced from Fig. 7 (right panel) for convenience. As with the other events, the transformer metric is the most performant, capturing more of the true flooding extents with fewer false positives.
Figure 13: PR curves for the Bangladesh flood scene. The transformer is the most performant model, achieving the largest area under the curve. The transformer also obtains the best $F_{1}$ score (starred for each model). The binary disturbance maps corresponding to the starred $F_{1}$ scores are shown in Fig. 12, third and fourth columns.